# QQESPM: A Quantitative and Qualitative Spatial Pattern Matching Algorithm

### v1

@@ 1 Introduction
While the SPM search methodology proves highly effective in scenarios necessitating distance constraints among queried POIs, it lacks the capability to address qualitative connectivity constraints between these entities. For instance, it cannot handle queries such as finding a school adjacent to a wooded area. To illustrate a more intricate search scenario, consider an individual seeking a rental space within a commercial building for their small business. In addition to this, they require the building to have an onsite gym and an adjacent green area. Furthermore, they need the building to be located within 1km from an elementary school for the convenience of their child’s enrollment and pickup. This complex scenario can be modeled using a spatial pattern graph that incorporates both quantitative (distance) and qualitative (connectivity) constraints, as shown in Figure 1 (B). However, existing SPM search algorithms are unable to handle such queries, requiring users to perform distance-based queries and manually sift through the results to find those meeting qualitative constraints.
Considering this challenge, this paper introduces a new type of POI group search: Qualitative and Quantitative Spatial Pattern Matching (QQ-SPM). The QQ-SPM query extends the conventional SPM query to encompass qualitative connectivity constraints between queried POIs. This approach enables the incorporation of qualitative requirements expressed through topological connectivity relations among the POI geometries. QQ-SPM thus provides a comprehensive solution that covers the entire spectrum of SPM search patterns while accommodating qualitative criteria specified by the user, enhancing the versatility of spatial pattern specification.
This work has the following key contributions:
@@ 2 Related Work
In this section we mention three of the main types of spatial keyword queries related to this work. The first type involves searching for POIs that meet specific keywords and are in close proximity to a designated center point for the search. For instance, the top-k spatial keyword search aims to identify geotextual objects (e.g., POIs) using a set of keywords and an initial search location. The goal is to locate the top-k closest POIs to the starting point while satisfying all search keywords. Studies in this field include those by [Cao et al. 2011, Hermoso et al. 2019, Zhang et al. 2013].
The second type of search focuses on minimizing distances between queried POIs. For instance, m-Closest spatial keyword search seeks groups of closely located POIs that collectively satisfy a user-defined set of $m$ keywords. Studies in this category include those conducted by [Choi et al. 2016, Choi et al. 2020, Guo et al. 2015]. However, the first two search types lack the capability to accommodate more intricate patterns, such as specifying a minimum distance between two returned POIs, which is essential when users want to avoid close proximity to certain types of POIs, like hospitals.
The third search type is the SPM search, which utilizes a graph-based spatial pattern. In this pattern, vertices store spatial keywords, and edges represent desired distance constraints. SPM search offers increased specificity by enabling users to impose both minimum and maximum distance restrictions between pairs of POIs. Studies in this category include works by [Fang et al. 2018a, Li et al. 2019, Fang et al. 2018b, Fang et al. 2019, Chen et al. 2019, Chen et al. 2022]. However, the SPM search lacks the capability to model qualitative restrictions, such as connectivity limitations.
In [Long et al. 2016], an efficient mechanism for indexing qualitative relations is proposed, aiming to reduce the time required for calculating the qualitative relation between two geometries. The core concept involves initially computing the qualitative relation between the Minimal Bounding Rectangles (MBRs) of the spatial objects. In cases where it is not possible to determine the topological relation between the geometries of the POIs solely based on the topological relation between their MBRs, their topological relation will be previously indexed. However, this approach primarily focuses on efficiently determining the qualitative relation between two existing geospatial objects within a dataset, rather than identifying the subset of objects satisfying a specific qualitative relation among numerous objects.
The concept of a spatial pattern defined by qualitative connectivity constraints is introduced in [Rafael 2021]. The work presents the Qualitative Spatial Pattern Search (QSPM) and an algorithm called the Topo-MSJ algorithm for addressing this type of search. However, being an early work in the realm of qualitative patterns, the author does not explore the potential of combining quantitative restrictions with qualitative ones in this search context.
@@ 3.1 IL-Quadtree
The Inverted Linear Quadtree (IL-Quadtree), a geotextual indexing structure introduced in [Zhang et al. 2016], serves as a fundamental component in the QQESPM algorithm. This index functions as a map, associating each unique spatial keyword in the dataset with its respective quadtree index structure. Each of these quadtrees contains a set of spatial objects (e.g., POIs) related to the specific keyword under consideration.
The bidimensional quadtree, as proposed in [Finkel and Bentley 1974], divides a two-dimensional spatial domain into four quadrants recursively. Each quadrant can be further subdivided into four subquadrants, and this subdivision is represented by a tree structure. Each rectangular subspace represents a node in the tree, and a node’s children correspond to its subquadrants. Subdivision occurs when the number of spatial entities (POIs) in a node exceeds a specified threshold, which can be adjusted during quadtree construction. Subspace division employs directional codes (00, 01, 10, and 11) to signify southwest, southeast, northwest, and northeast quadrants, respectively. Concatenating these codes recursively provides a unique identifier for each node, indicating its position in the quadtree hierarchy. Figure 2 illustrates the spatial partitioning in the quadtree and its associated tree structure. The IL-Quadtree’s architecture efficiently retrieves geotextual objects during geotextual queries, as indicated in [Zhang et al. 2016] and [Chen et al. 2019]. The QQESPM algorithm relies on the IL-Quadtree indexing method to perform queries.
@@ 3.2 DE-9IM
A foundational model for computing the topological connectivity relation between two-dimensional geometries is the Dimensionally Extended Nine-Intersection Model (DE-9IM) [Egenhofer and Herring 1990, Clementini et al. 1993, Clementini et al. 1994]. This representation provides a structured framework for formally defining spatial predicates that describe the connectivity between POIs. DE-9IM can represent topological relations such as “equals”, “touches” and “contains”.
This topological relation model utilizes a 3x3 matrix to represent the topological relation between two distinct geometries, denoted as A and B. The matrix elements represent intersections across the interior, boundary, and exterior components of these geometries. Each matrix configuration corresponds to a possible topological relation. A simple description for some topological relations can be found in Table 1.
Table 1: Topological Predicates [Strobl 2008]
@@
Equals
The Geometries are topologically equal
Disjoint
The Geometries have no point in common
Intersects
The Geometries have at least one point in common (the inverse of Disjoint)
Touches
The Geometries have at least one boundary point in common, but no interior points
Partially Overlaps
The Geometries share some but not all points in common, and the intersection has the same dimension as the Geometries themselves
Within
Geometry A lies in the interior of Geometry B
Contains
Geometry B lies in the interior of Geometry A (the inverse of Within)
@@ 4 Problem Formalization
Edges with the distance interval information are called quantitative edges, and edges with the connectivity predicate are called qualitative edges. Edges may or may not be simultaneously quantitave and qualitative. If a quantitative edge has the mutual inclusion sign, it is called an inclusive edge, otherwise, it is called an exclusive edge. Also note that, since the relation “covered by” is the inverse of “covers”, it could be discarded, but once edges are directional, i.e., have specific starting and ending vertices, we keep the relation “covered by”.
Notice that the attributes of an edge for the QQ-SPM query is a generalization of the attributes of an edge for the SPM query allowing qualitative connectivity constraints. In this way, the spatial pattern definition for the QQ-SPM query is also a generalization of the spatial pattern definition for the SPM query.
@@
The proof for the quantitative restrictions of the edges is provided in [Chen et al. 2019]. Regarding the additional proposed criterion to qualify as a qq-n-match, which is related to the connectivity constraint of the edge, it’s important to note that if the node pair $(n_{i},n_{j})$ constitutes a qq-n-match for an edge with a qualitative constraint other than disjoint, they will possess intersecting bounding boxes. Since their father nodes encompass them, they too will be intersecting, thus ensuring that the condition for a node match persists for the father nodes.
@@ 5 QQESPM algorithm
1 $L=max(depth(ILQ_{i}),1\leq i\leq n)$ for $level=1$ to $L$ do
2       derive the order of computing qq-n-matches for this level for each edge $e$ do
@@ 1 $L=max(depth(ILQ_{i}),1\leq i\leq n)$ for $level=1$ to $L$ do
4
5derive the order of computing qq-e-matches identify skip-edges for each non-skip edge $e$ do
@@
derive the order of joining qq-e-matches $\psi\leftarrow join\_qq\_e\_matches()$ return $\psi$
@@
This section presents the QQESPM algorithm, designed to handle QQ-SPM queries. QQESPM considers six possible topological relations between POIs, namely “equals”, “touches”, “covers”, “covered by”, “partially overlaps”, and “disjoint”. The overview of the search procedure is shown in Algorithm 1 (QQESPM), which delineates the high-level sequential steps for query execution, with an emphasis on achieving efficient execution by using the qq-n-matches concept. It iteratively operates at the depth levels in the quadtrees of keywords in the search pattern, examining qq-n-matches for each edge at the current depth level by evaluating child node pairs from the previous depth level (according to Lemma 1). Upon reaching the final depth level of the quadtrees, it tests the pairs of objects within the last level’s qq-n-matches to identify qq-e-matches. The qq-e-matches of each edge are then joined to produce solutions (matches) for the spatial pattern.
Within the context of a given edge $e(v_{i},v_{j})$, the QQESPM algorithm systematically searches for its qq-n-matches in a tiered manner, starting from the root nodes of the $ILQ_{i}$ and $ILQ_{j}$ quadtrees. The initial qq-n-match for edge $e$ arises from the pair of root nodes, specifically (root of $ILQ_{i}$, root of $ILQ_{j}$), exclusively at level 0. The process then progresses to examine pairs of nodes $(n_{i},n_{j})$, where $n_{i}$ is a child of the root of $ILQ_{i}$ and $n_{j}$ is a child of the root of $ILQ_{j}$, following Lemma 1. This step identifies the qq-n-matches for $e$ at the 1st level. The algorithm continues this exploration iteratively by inspecting the children of these nodes, deducing qq-n-matches for $e$ at the 2nd level and subsequent levels. This iterative traversal persists until the maximum depth of the quadtrees is reached. The final level’s qq-n-matches are retained to subsequently derive the qq-e-matches for the edges.
At each level, the algorithm employs a reordering strategy for the edges list, giving priority to edges with fewer qq-n-matches from the previous level, guided by the rationale that such edges are more likely to yield fewer qq-e-matches. This strategic ordering accelerates computation by swiftly eliminating unsuitable nodes early, as proposed by [Chen et al. 2019].
After calculating qq-n-matches for all edges at the final level (the maximum depth of the quadtrees), the algorithm evaluates POI pairs within each edge’s qq-n-matches to determine the qq-e-matches. Before computing qq-e-matches for any edge, the algorithm checks if the terminal vertices of the edge have a restricted set of candidate objects. This set is obtained from qq-e-matches of edges with shared vertices. This strategy serves as a pre-joining mechanism avoiding redundant pair assessments. Also, the calculation of qq-e-matches is not necessary for some mutually inclusive edges whose extreme vertices are shared with other edges whose qq-e-matches will be computed, so that for these edges, called skip edges, the verification of constraints occurs at the joining stage, which compares the qq-e-matches of edges with shared vertices, and eliminates the non-matching ones.
The structural framework and strategic heuristics of ESPM are replicated within the QQESPM algorithm. The divergence lies in the criteria for qq-n-match and qq-e-match identification, as defined in Definitions 2 and 4, which will occur in lines 5 and 9. These divergent criteria is sufficient to promote distinct computations at each level, as qq-n-matches are computed level by level from the root to the maximum depth level of the keyword’s quadtrees.
@@ 6.1 Experiments Description
The ESPM algorithm was implemented in Python, following the description provided in [Chen et al. 2019]. This implementation was further adapted to include the qq-n-matches and qq-e-matches verification stages to accommodate qualitative connectivity constraints, resulting in the initial implementation of the QQESPM algorithm [^3] The code for this implementation can be found in https://zenodo.org/records/10085300. Additionally, a more straightforward approach, referred to as QQ-simple, was also implemented to be compared with QQESPM. This approach checks the connectivity constraints only at the final step of ESPM, employing a filtering mechanism to exclude solutions that do not meet the connectivity requirements. Subsequently, we conducted a comparative performance analysis of these two QQ-SPM solutions.
Experiments were executed on a machine equipped with Intel Core i7-12700F CPU 4.90 GHz, coupled with 32GB of memory, operating on the Ubuntu OS. The computational load was carried out on a single CPU core.
We used a dataset comprising 33,877 POIs, extracted from OpenStreetMap [^4] https://www.openstreetmap.org/ filtered by the following bounding box: {min_lat: -8.3610, min_long:-38.8559, max_lat: -5.9275, max_long: -34.7415}, thereby predominantly spanning the Paraíba state, Brazil. The dataset comprises the tags ‘amenity’, ‘shop’, and ‘tourism’, containing 315 distinct keywords.
In an effort to construct resource-intensive search spatial patterns, mirroring real-world conditions, the 20 most frequent keywords were identified and selected from the tags ‘amenity’, ‘shop’, and ‘tourism’, amounting to a cumulative set of 60 keywords to compose the search patterns.
The experiments encompassed 12 distinct graph architectures for spatial patterns, following [Chen et al. 2019]. These structural patterns can be visualized in Figure 3. For each of these architectures, 5 distinct spatial patterns were generated with randomly selected keywords, totalizing 60 spatial patterns.
The dataset was randomly shuffled and divided into segments representing $20\%$, $40\%$, $60\%$, $80\%$, and $100\%$ of the total POI count. For each of these dataset subsets, searches were conducted five times for each of the 60 spatial patterns generated, for both algorithms QQESPM and QQ-simple. Thus, the total number of executions was 3,000, and each of the two algorithms answered 1,500 queries.
For the purpose of this study, a simplified convention was adopted, by using the Euclidean distance to measure the distance between POIs coordinates (longitude and latitude). Note that it differs from the distance in kilometers. To construct the search patterns, the parameter $l_{ij}$, representing the minimum inter-POI distance, was randomly drawn from the interval $[0,0.005]$ (equivalent to approximately 00 to $550m$), while the parameter $u_{ij}$, representing the maximum inter-POI distance, was randomized within the range $[l_{ij},l_{ij}+0.02]$ (reaching up to $2.7km$ approximately). The connectivity relations were introduced in the edges randomly from the set {“equals” , “touches”, “covers”, “covered by”, “partially overlaps”, “disjoint”}. Each edge had a probability of $50\%$ of receiving a connectivity relation constraint.

### v2

@@ 1 Introduction
While the SPM search methodology proves highly effective in scenarios necessitating distance constraints among queried POIs, it lacks the capability to address qualitative connectivity constraints between these entities. For instance, it cannot handle queries such as finding a school adjacent to a wooded area. To illustrate a more intricate search scenario, consider an individual seeking a rental space within a commercial building for their small business. In addition to this, they require the building to have an onsite gym and an adjacent green area. Furthermore, they need the building to be located within 1km from an elementary school for the convenience of the enrollment and pickup of their child. This complex scenario can be modeled using a spatial pattern graph that incorporates both quantitative (distance) and qualitative (connectivity) constraints, as shown in Figure 1 (B). However, existing SPM search algorithms are unable to handle such queries, requiring users to perform distance-based queries and manually sift through the results to find those meeting qualitative constraints.
@@
Considering this challenge, this paper introduces a new type of POI group search: Qualitative and Quantitative Spatial Pattern Matching (QQ-SPM). The QQ-SPM query extends the conventional SPM query to encompass qualitative connectivity constraints between queried POIs. This approach enables the incorporation of qualitative requirements expressed through topological connectivity relations among the POI geometries. QQ-SPM thus provides a comprehensive solution that covers the entire spectrum of SPM search patterns while accommodating qualitative criteria specified by the user, enhancing the versatility of spatial pattern specification. This work has the following key contributions:
@@ 2 Related Work
In this section we mention three of the main types of spatial keyword queries related to this work. The first type involves searching for POIs that meet specific keywords and are in close proximity to a designated center point for the search. For instance, the top-k spatial keyword search aims to identify geotextual objects (e.g., POIs) using a set of keywords and an initial search location. The goal is to locate the top-k closest POIs to the starting point while satisfying all search keywords. Studies in this field include those by [Cao et al. 2011, Hermoso et al. 2019, Zhang et al. 2013]. The second type of search focuses on minimizing distances between queried POIs. For instance, m-Closest spatial keyword search seeks groups of closely located POIs that collectively satisfy a user-defined set of $m$ keywords. Studies in this category include those conducted by [Choi et al. 2016, Choi et al. 2020, Guo et al. 2015]. However, the first two search types lack the capability to accommodate more intricate patterns, such as specifying a minimum distance between two returned POIs, which is essential when users want to avoid close proximity to certain types of POIs, like hospitals. The third search type is the SPM search, which utilizes a graph-based spatial pattern. In this pattern, vertices store spatial keywords, and edges represent desired distance constraints. SPM search offers increased specificity by enabling users to impose both minimum and maximum distance restrictions between pairs of POIs. Studies in this category include works by [Fang et al. 2018a, Li et al. 2019, Fang et al. 2018b, Fang et al. 2019, Chen et al. 2019, Chen et al. 2022]. However, the SPM search lacks the capability to model qualitative restrictions, such as connectivity limitations. In [Long et al. 2016], an efficient mechanism for indexing qualitative relations is proposed, aiming to reduce the time required for calculating the qualitative relation between two geometries. The core concept involves initially computing the qualitative relation between the Minimal Bounding Rectangles (MBRs) of the spatial objects. In cases where it is not possible to determine the topological relation between the geometries of the POIs solely based on the topological relation between their MBRs, their topological relation will be previously indexed. However, this approach primarily focuses on efficiently determining the qualitative relation between two existing geospatial objects within a dataset, rather than identifying the subset of objects satisfying a specific qualitative relation among numerous objects. The concept of a spatial pattern defined by qualitative connectivity constraints is introduced in [Rafael 2021]. The work presents the Qualitative Spatial Pattern Search (QSPM) and an algorithm called the Topo-MSJ algorithm for addressing this type of search. However, being an early work in the realm of qualitative patterns, the author does not explore the potential of combining quantitative restrictions with qualitative ones in this search context.
@@ 3.1 IL-Quadtree
The Inverted Linear Quadtree (IL-Quadtree), a geotextual indexing structure introduced in [Zhang et al. 2016], serves as a fundamental component in the QQESPM algorithm. This index functions as a map, associating each unique spatial keyword in the dataset with its respective quadtree index structure. Each of these quadtrees contains a set of spatial objects (e.g., POIs) related to the specific keyword under consideration. The bidimensional quadtree, as proposed in [Finkel and Bentley 1974], divides a two-dimensional spatial domain into four quadrants recursively. Each quadrant can be further subdivided into four subquadrants, and this subdivision is represented by a tree structure. Each rectangular subspace represents a node in the tree, and a node’s children correspond to its subquadrants. Subdivision occurs when the number of spatial entities (POIs) in a node exceeds a specified threshold, which can be adjusted during quadtree construction. Subspace division employs directional codes (00, 01, 10, and 11) to signify southwest, southeast, northwest, and northeast quadrants, respectively. Concatenating these codes recursively provides a unique identifier for each node, indicating its position in the quadtree hierarchy. Figure 2 illustrates the spatial partitioning in the quadtree and its associated tree structure. The IL-Quadtree’s architecture efficiently retrieves geotextual objects during geotextual queries, as indicated in [Zhang et al. 2016] and [Chen et al. 2019]. The QQESPM algorithm relies on the IL-Quadtree indexing method to perform queries.
@@ 3.2 DE-9IM
A foundational model for computing the topological connectivity relation between two-dimensional geometries is the Dimensionally Extended Nine-Intersection Model (DE-9IM) [Egenhofer and Herring 1990, Clementini et al. 1993, Clementini et al. 1994]. This representation provides a structured framework for formally defining spatial predicates that describe the connectivity between POIs. DE-9IM can represent topological relations such as “equals”, “touches” and “contains”. This topological relation model utilizes a 3x3 matrix to represent the topological relation between two distinct geometries, denoted as A and B. The matrix elements represent intersections across the interior, boundary, and exterior components of these geometries. Each matrix configuration corresponds to a possible topological relation. A simple description for some topological relations from [Strobl 2008] can be found in Table 1.
Table 1: Topological Predicates
@@
Equals The Geometries are topologically equal
Disjoint The Geometries have no point in common
Intersects The Geometries have at least one point in common (the inverse of Disjoint)
Touches The Geometries have at least one boundary point in common, but no interior points
Partially Overlaps The Geometries share some but not all points in common, and the intersection has the same dimension as the Geometries themselves
Within Geometry A lies in the interior of Geometry B
Contains Geometry B lies in the interior of Geometry A (the inverse of Within)
@@ 4 Problem Formalization
Edges with the distance interval information are called quantitative edges, and edges with the connectivity predicate are called qualitative edges. Edges may or may not be simultaneously quantitave and qualitative. If a quantitative edge has the mutual inclusion sign, it is called an inclusive edge, otherwise, it is called an exclusive edge. Also note that, since the relation “covered by” is the inverse of “covers”, it could be discarded, but once edges are directional, i.e., have specific starting and ending vertices, we keep the relation “covered by”. Notice that the attributes of an edge for the QQ-SPM query is a generalization of the attributes of an edge for the SPM query allowing qualitative connectivity constraints. In this way, the spatial pattern definition for the QQ-SPM query is also a generalization of the spatial pattern definition for the SPM query.
@@
The proof for the quantitative restrictions of the edges is provided in [Chen et al. 2019]. Regarding the additional proposed criterion to qualify as a qq-n-match, which is related to the connectivity constraint of the edge, it’s important to note that if the node pair $(n_{i},n_{j})$ constitutes a qq-n-match for an edge with a qualitative constraint other than disjoint, they will possess intersecting bounding boxes. Since their father nodes encompass them, they too will be intersecting, thus ensuring that the condition for a node match persists for the father nodes. ∎
@@ 5 QQESPM algorithm
1 $L=max(depth(ILQ_{i}),1\leq i\leq n)$
2 for $level=1$ to $L$ do
3       derive the order of computing qq-n-matches for this level
4
5      for each edge $e$ do
@@ 1 $L=max(depth(ILQ_{i}),1\leq i\leq n)$
7
8
9derive the order of computing qq-e-matches
10 identify skip-edges
11 for each non-skip edge $e$ do
@@
13
14derive the order of joining qq-e-matches
15 $\psi\leftarrow join\_qq\_e\_matches()$
16 return $\psi$
@@ 15 $\psi\leftarrow join\_qq\_e\_matches()$
This section presents the QQESPM algorithm, designed to handle QQ-SPM queries. QQESPM considers six possible topological relations between POIs, namely “equals”, “touches”, “covers”, “covered by”, “partially overlaps”, and “disjoint”. The overview of the search procedure is shown in Algorithm 1 (QQESPM), which delineates the high-level sequential steps for query execution, with an emphasis on achieving efficient execution by using the qq-n-matches concept. It iteratively operates at the depth levels in the quadtrees of keywords in the search pattern, examining qq-n-matches for each edge at the current depth level by evaluating child node pairs from the previous depth level (according to Lemma 1). Upon reaching the final depth level of the quadtrees, it tests the pairs of objects within the last level’s qq-n-matches to identify qq-e-matches. The qq-e-matches of each edge are then joined to produce solutions (matches) for the spatial pattern. Within the context of a given edge $e(v_{i},v_{j})$, the QQESPM algorithm systematically searches for its qq-n-matches in a tiered manner, starting from the root nodes of the $ILQ_{i}$ and $ILQ_{j}$ quadtrees. The initial qq-n-match for edge $e$ arises from the pair of root nodes, specifically (root of $ILQ_{i}$, root of $ILQ_{j}$), exclusively at level 0. The process then progresses to examine pairs of nodes $(n_{i},n_{j})$, where $n_{i}$ is a child of the root of $ILQ_{i}$ and $n_{j}$ is a child of the root of $ILQ_{j}$, following Lemma 1. This step identifies the qq-n-matches for $e$ at the 1st level. The algorithm continues this exploration iteratively by inspecting the children of these nodes, deducing qq-n-matches for $e$ at the 2nd level and subsequent levels. This iterative traversal persists until the maximum depth of the quadtrees is reached. The final level’s qq-n-matches are retained to subsequently derive the qq-e-matches for the edges. At each level, the algorithm employs a reordering strategy for the edges list, giving priority to edges with fewer qq-n-matches from the previous level, guided by the rationale that such edges are more likely to yield fewer qq-e-matches. This strategic ordering accelerates computation by swiftly eliminating unsuitable nodes early, as proposed by [Chen et al. 2019]. After calculating qq-n-matches for all edges at the final level (the maximum depth of the quadtrees), the algorithm evaluates POI pairs within each edge’s qq-n-matches to determine the qq-e-matches. Before computing qq-e-matches for any edge, the algorithm checks if the terminal vertices of the edge have a restricted set of candidate objects. This set is obtained from qq-e-matches of edges with shared vertices. This strategy serves as a pre-joining mechanism avoiding redundant pair assessments. Also, the calculation of qq-e-matches is not necessary for some mutually inclusive edges whose extreme vertices are shared with other edges whose qq-e-matches will be computed, so that for these edges, called skip edges, the verification of constraints occurs at the joining stage, which compares the qq-e-matches of edges with shared vertices, and eliminates the non-matching ones. The structural framework and strategic heuristics of ESPM are replicated within the QQESPM algorithm. The divergence lies in the criteria for qq-n-match and qq-e-match identification, as defined in Definitions 2 and 4, which will occur in lines 5 and 9. These divergent criteria is sufficient to promote distinct computations at each level, as qq-n-matches are computed level by level from the root to the maximum depth level of the keyword’s quadtrees.
@@ 6.1 Experiments Description
The ESPM algorithm was implemented in Python, following the description provided in [Chen et al. 2019]. This implementation was further adapted to include the qq-n-matches and qq-e-matches verification stages to accommodate qualitative connectivity constraints, resulting in the initial implementation of the QQESPM algorithm [^3] The code for this implementation can be found in https://zenodo.org/records/10085300. Additionally, a more straightforward approach, referred to as QQ-simple, was also implemented to be compared with QQESPM. This approach checks the connectivity constraints only at the final step of ESPM, employing a filtering mechanism to exclude solutions that do not meet the connectivity requirements. Subsequently, we conducted a comparative performance analysis of these two QQ-SPM solutions. Experiments were executed on a machine equipped with Intel Core i7-12700F CPU 4.90 GHz, coupled with 32GB of memory, operating on the Ubuntu OS. The computational load was carried out on a single CPU core. We used a dataset comprising 33,877 POIs, extracted from OpenStreetMap [^4] https://www.openstreetmap.org/ filtered by the following bounding box: {min_lat: -8.3610, min_long:-38.8559, max_lat: -5.9275, max_long: -34.7415}, thereby predominantly spanning the Paraíba state, Brazil. The dataset comprises the tags ‘amenity’, ‘shop’, and ‘tourism’, containing 315 distinct keywords. In an effort to construct resource-intensive search spatial patterns, mirroring real-world conditions, the 20 most frequent keywords were identified and selected from the tags ‘amenity’, ‘shop’, and ‘tourism’, amounting to a cumulative set of 60 keywords to compose the search patterns. The experiments encompassed 12 distinct graph architectures for spatial patterns, following [Chen et al. 2019]. These structural patterns can be visualized in Figure 3. For each of these architectures, 5 distinct spatial patterns were generated with randomly selected keywords, totalizing 60 spatial patterns. The dataset was randomly shuffled and divided into segments representing $20\%$, $40\%$, $60\%$, $80\%$, and $100\%$ of the total POI count. For each of these dataset subsets, searches were conducted five times for each of the 60 spatial patterns generated, for both algorithms QQESPM and QQ-simple. Thus, the total number of executions was 3,000, and each of the two algorithms answered 1,500 queries. For the purpose of this study, a simplified convention was adopted, by using the Euclidean distance to measure the distance between POIs coordinates (longitude and latitude). Note that it differs from the distance in kilometers. To construct the search patterns, the parameter $l_{ij}$, representing the minimum inter-POI distance, was randomly drawn from the interval $[0,0.005]$ (equivalent to approximately 00 to $550m$), while the parameter $u_{ij}$, representing the maximum inter-POI distance, was randomized within the range $[l_{ij},l_{ij}+0.02]$ (reaching up to $2.7km$ approximately). The connectivity relations were introduced in the edges randomly from the set {“equals” , “touches”, “covers”, “covered by”, “partially overlaps”, “disjoint”}. Each edge had a probability of $50\%$ of receiving a connectivity relation constraint.